Crcvit -(20) or Ahe given Qr And S and G14) Volts Gle) = 100 cosine t
Answer #1

given valt) = 100 Cosinect) volts. Radian frequency 60 = l rod/sec. In phasor domain Vg = 100 (0% (Vg)rms = 100 100 Inductivej22 ro Reo=22 16 100 do = 109 10° Reqt XL 2+j2 a=251-45 A The complex power supplied by sare S= Vg a* = 10060° * (25145) = 10